

********************************************************************
UARS WEEKLY STATUS REPORT - 01 July 2005
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
67% of each day's data amounting to about 711 minutes of solar
observing time per day.
***********
SUSIM
***********
During this past week, SUSIM has successfully gathered its daily mid
(1.1 nm) resolution, MgII, and continuous monitoring scans for time
periods with UARS telemetry or tape recording. An offset dwell and
O2 Occultation were also performed.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005059 has been processed and is waiting for archival
at the DAAC. HEPSA data for 2005060 to 2005069 is currently being processed.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005059 has been processed and is waiting for DAAC archival. MEPS data for
2005060 to 2005069 is currently being processed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 18 June 2005 (Orbit 75362, GMT Day 169) through
24 June 2005 (Orbit 75466, GMT Day 175).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 170/2009 75389 DAY-ONLY TO FULL-TIME
PEM 172/1604 75417 FULL-TIME TO DAY-ONLY
The beta angle ranged between the angles of 77.8 degrees to 68.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 1 degree on 10 July 2005 (DOY 191).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 2.53 to 2.20 degrees C. Additional
battery performance data for each battery on 17 June 2005 and 24 June
2005 are provided in Appendix D.
The clock error ranged from 10.0 to -10.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
PEM = 02
SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 05
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Five (5) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to closely monitor the status of the
batteries. Over the last several days UARS has been flying through
its bi-annual high beta period. During this time the day portion of
the orbit is extended and more power is available to support payloads.
However, due to HALOE's temperature constraint, it must be powered off
the keep it from over heating. Therefore given the increased power
available, the battery charge level was lowered to VT-3 on 18 June
2005 on the 169/1113 event. As the beta angle began to move back down
from its peak of 80 degrees, the charge level was raised back up to VT-
4 on 21 June 2005 on the 172/2226 event. Throughout this period the
batteries were able to support SUSIM and SOLSTICE fulltime.
The UARS OBC clock rollover effort continued this week with ATC
commanding added to the uplink procedure. This addition was tested on
the simulator and the uplink procedure is now fairly complete. There
may be some additional tweaking over the course of the next few weeks
but no major changes are expected. Also, the ground system transition
plan is being worked to handle updating the ground system with the new
epoch year. A status on the progress of the Eclipse development team
will be received on 28 June 2005.
MEETINGS
The FOT attended a telecon with the Eclipse development team on 21
June 2005 to discuss the patch needed for the ground system to accept
the new onboard epoch year.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 10 July 2005 (DOY 191). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 24 June 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
75369 169/1113 05-193 MPS TO VT L-3 (DUE TO HIGH BETA)
75375 169/2056 05-194 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
75389 170/2009 05-195 PEM DAY ONLY TO FULL (POWER LOAD)
75410 172/0444 05-196 OBC/CLK CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
75417 172/1604 05-197 PEM FULLTIME OPS TO DAY ONLY
75421 172/2226 05-198 MPS VT L-3 TO VT L-4
75434 173/1958 05-197 PEM PEM DAY ONLY CMD
75436 173/2222 05-199 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
75462 175/1526 05-200 SUSIM EXCEPTION CLEAR
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 24 June 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 24 June 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
75375 169/2059 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
75410 172/0452 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
75436 173/2222 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 24 June 2005
17 June 2005 - GMT Day 168 Beta = 75.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 00.0/ -22.4 -3.05/-4.62 +00.0/+00.0 05.8 0.00 0.0
2 00.0/ -89.6 +3.87/+3.23 +18.4/-20.0 27.0 1.05 15.4
3 00.0/-700.0 +1.34/+0.09 +09.6/-08.8 26.9 1.08 06.9
24 June 2005 - GMT Day 175 Beta = 68.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 11.2/ -22.4 -3.05/-3.99 +00.0/+00.0 05.8 0.00 0.0
2 -89.6/-201.6 +4.19/+3.23 +20.8/-16.4 26.2 1.03 16.5
3 22.4/-700.0 +1.66/+1.03 +14.4/-10.0 26.2 1.06 10.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
75369 169/1113 LV4 -> LV3
75421 172/2226 LV3 -> LV4
APPENDIX E
UARS ANOMALY REPORTS
Week ending 24 June 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-030 FOT 75392 171/0047 30832 Generic Late Acquisition
05-030 FOT 75410 172/0444 30832 Generic Late Acquisition
05-030 FOT 75414 172/1102 30832 Generic Late Acquisition
05-030 FOT 75439 174/0256 30832 Generic Late Acquisition
05-030 FOT 75446 174/1417 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-030 FOT See List June 05 30832 Generic Late Acquisition